OpenStudy (anonymous):

1+1+1/3 There is two answers to this :) Just wondering who has an open mind.

hartnn (hartnn):

7/3 and 1 ?

hartnn (hartnn):

naaah, not 1....for 1 , brackets are essential (1+1+1)/3

OpenStudy (anonymous):

it has to be just 7/3 =2 1/3 mixed fraction

OpenStudy (anonymous):

or as hartnn has pointed out

hartnn (hartnn):

or do u mean 2 equivalent answers like 7/3 and \(2\frac{1}{3}\)

OpenStudy (anonymous):

There no right or wrong since I didn't point out a method I'm just wondering what you all can come up with :)

OpenStudy (anonymous):

\[1+1+1/3=1+1+\frac{1}{3}=2+\frac{1}{3} \] Multiplication has precedence over addition.

OpenStudy (anonymous):

There are not two possible interpretations. An open mind would understand there is only one. 1+1+1/3 = 2 1/3 (1+1) + 1/3 = 2 1/3 (1+1+1)/3 = 1

OpenStudy (anonymous):

Sorry. My bad. Division has precedence over addition.

OpenStudy (shubhamsrg):

if we are allowed to put brackets anywhere,, how about (1+1+1)/3 1+ (1+1)/3 and 1+1 +(1/3)

OpenStudy (anonymous):

1+1 =2 2+1/3 =(2(3) +1)/3 =(6+1)/3 =7/3 you can write this as mixed fraction as well which is \[=2\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

Placing parentheses willy nilly can change the original expression's final form and the resulting expression's value. 1+1+1/3 = 2 and 1/3 or 7/3 and that's that.

OpenStudy (anonymous):

1+1+1/3 is only equal to 2.333333333333334 and that's that. lol